# InternVLA-A1: Unifying Understanding, Generation and Action for Robotic Manipulation

### v1

@@ 5.5 Ablation Studies
Effectiveness of Pre-training. As shown in Figure 7, we assess the impact of pre-training by comparing our model with a version trained from scratch. The removal of the pre-training stage resulted in an overall performance drop of 51.6%, with the average success rate falling from 77.0% to 25.4%. In severe cases, the baseline failed completely, while our model retained high proficiency. This finding suggests that pre-training acts as a crucial inductive prior, effectively stabilizing the optimization process and allowing the model to inherit foundational manipulation skills. As a result, our model achieves a significantly higher success rate on downstream tasks by building upon the robust capabilities acquired during pre-training.
@@
Effectiveness of Generation Expert. To evaluate the contribution of the generation expert, we conduct a key ablation study by comparing InternVLA-A1 (3B) with that without the generation expert. The experimental results are shown in Figure˜8. The results demonstrate that InternVLA-A1 (3B) outperforms the ablated version (without the generation expert) in 11 out of 12 real-world tasks, achieving performance gains ranging from 6.7% to 53.3%. Notably, removing the generation expert significantly reduces the average success rate from 77.0% to 57.6%. This ablation study validates the superiority of the proposed generation expert and the unified architecture integrating understanding, generation, and action.
Impact of Pre-training Datasets. As shown in Figure 9, we evaluate the efficacy of different pre-training data sources: the large-scale real-world dataset Agibot-World, our high-fidelity synthetic dataset InternData-A1, and a mixed strategy combining both. First, InternData-A1 demonstrates exceptional transfer capabilities. Despite being synthetic, the model pre-trained solely on InternData-A1 matches or surpasses the real-data-only baseline, confirming InternData-A1’s high fidelity for supporting pre-training. Second, InternData-A1 dominates in simulation benchmarks. On the RoboTwin simulation evaluation, the pure InternData-A1 model achieves the highest performance, outperforming even the mixed pre-training setting. We attribute this to two factors: (1) the minimal visual domain gap between the training and evaluation environments (Sim-to-Sim), and (2) the diverse range of heterogeneous embodiments within InternData-A1, which provides better adaptability to the target robot compared to the specific embodiment found in Agibot-World. Finally, mixed pre-training synergizes complementary strengths. While adding real-world data causes a slight performance decline in simulation compared to pure synthetic training, the mixed strategy yields substantial gains in real-world scenarios, particularly in fine-grained manipulation tasks such as Sort Parts, where we observe a 20% improvement. This confirms that our hybrid approach effectively synergizes the scale and diversity of simulation with the physical realism of real-world data.
@@ 6 Conclusion and Limitations
In this work, we presented InternVLA-A1, a VLA model with enhanced robustness against environmental dynamics. We achieved this through dual advancements in model architecture and data paradigms. Architecturally, InternVLA-A1 employs a Mixture-of-Transformers (MoT) design to unify semantic understanding, visual foresight, and action prediction, effectively synergizing high-level reasoning with low-level dynamics. Complementing this, our hybrid synthetic-real pre-training strategy combines the scene diversity of simulation with the physical fidelity of real-world data. Extensive evaluations demonstrate that InternVLA-A1 significantly outperforms leading baselines such as and GR00T, exhibiting exceptional robustness, particularly in highly dynamic scenarios. Furthermore, ablation studies validate the critical contributions of the proposed generation expert, the unified architecture, and the complementary efficacy of our hybrid data strategy.
Limitations. Despite these advancements, two primary limitations remain. First, the understanding expert currently lacks joint training with multimodal VQA datasets, leading to a degradation in general semantic reasoning and complex instruction-following capabilities. Second, to ensure efficient inference for the visual foresight module, we compromised the fidelity of image prediction, which limits the granularity of the generated future frames. Future work will address these two limitations.
Appendix A Appendix
@@
Junhao Cai1, Yang Li1, Jiangmiao Pang1†, Zherui Qiu1, Yang Tian1, Jia Zeng1†
@@
Zetao Cai1, Jiafei Cao1, Yilun Chen1, Zeyu He1, Lei Jiang2, Hang Li1, Hengjie Li1, Yufei Liu2, Yanan Lu1, Qi Lv1, Haoxiang Ma1, Yu Qiao1, Yanqing Shen1, Xu Shi1, Bolun Wang1, Hanqing Wang1, Jiaheng Wang1, Tai Wang1, Xueyuan Wei1, Chao Wu1, Yiman Xie1, Boyang Xing2, Yuqiang Yang1, Yuyin Yang1, Qiaojun Yu1, Feng Yuan1, Jingjing Zhang1, Shenghan Zhang1, Shi Zhang1, Zhuoma Zhaxi1, Bowen Zhou1, Yuanzhen Zhou1, Yunsong Zhou1, Hongrui Zhu1, Yangkun Zhu1, Yuchen Zhu2

### v2

@@ 5.5 Ablation Studies
Impact of Pre-training. As shown in Figure 8, we assess the impact of pre-training by comparing our model with a version trained from scratch. The removal of the pre-training stage resulted in an overall performance drop of 51.6%, with the average success rate falling from 77.0% to 25.4%. In severe cases, the baseline failed completely, while the pre-trained model retained high proficiency. This finding suggests that pre-training acts as a crucial inductive prior.
Impact of Pre-training Datasets. As shown in Table˜5, we evaluate the efficacy of different pre-training data sources. Training on simulation data alone already yields strong results on RoboTwin 2.0 (88.3%/88.5% on Easy/Hard), but generalizes less well to real-world tasks (53.3% on Place flower and 33.3% on Sort parts). Incorporating human videos improves simulation performance (up to 89.4%/89.3%) and slightly benefits real-world generalization (specifically improving Sort parts to 40.0%). Most notably, jointly pre-training on heterogeneous data sources (human videos, synthetic data, and real-world demonstrations) achieves the best overall performance, and substantially enhances real-world manipulation performance (reaching 60.0% on Place flower and 53.3% on Sort parts). This demonstrates the effectiveness of our joint training strategy.
| Pre-training dataset | RoboTwin 2.0 | Real-world tasks | ||
| Easy | Hard | Place flower | Sort parts | |
| Sim. only | 88.3 | 88.5 | 53.3 | 33.3 |
| Sim. + Human | 89.4 | 89.3 | 53.3 | 40.0 |
| Sim. + Real + Human | 89.4 | 89.6 | 60.0 | 53.3 |
@@
Impact of Generation Expert. To evaluate the contribution of the generation expert, we conduct a key ablation study by comparing InternVLA-A1 (3B) with the variant without the generation expert. The experimental results are presented in Figure˜9. The results demonstrate that InternVLA-A1 (3B) outperforms the ablated version (without the generation expert) in 11 out of 12 real-world tasks, achieving performance gains ranging from 6.7% to 53.3%. Notably, removing the generation expert significantly reduces the average success rate from 77.0% to 57.6%, with the most substantial decline observed in dynamic manipulation tasks ( Express Sorting and In-motion Ingredient Picking). This ablation study validates the superiority of the proposed generation expert and the unified architecture integrating understanding, generation, and action.
5.6 Visualization
The head-view future predictions generated by InternVLA-A1 are illustrated in Figure˜10. As shown in the samples, the predictions sacrifice some high-frequency visual details. This is a deliberate design choice to prioritize inference efficiency over pixel-level granularity. Despite this, the model accurately captures the essential motion trends and dynamics. We argue that when integrating visual foresight generation module into manipulation policies, high-frequency detail or visual clarity is secondary; what matters most is that the latent features encapsulate sufficient instructive information to guide action execution.
@@ 6 Conclusion and Limitations
In this work, we presented InternVLA-A1, a unified framework that integrates scene understanding, visual-foresight generation, and action execution through a Mixture-of-Transformers (MoT) design. This architecture couples semantic reasoning with dynamics prediction, and enables effective joint training on heterogeneous data sources (human videos, synthetic data, and real-world demonstrations). As a result, InternVLA-A1 achieves consistent robustness across static manipulation, dynamic manipulation, and simulation benchmarks, with particularly strong gains in highly dynamic scenarios.
Limitations. Despite these advancements, two primary limitations remain. First, the understanding expert is not jointly trained with large-scale multimodal VQA data, which weakens general semantic reasoning and complex instruction following. Second, to ensure efficient inference for the visual foresight generation module, we compromised the fidelity of image prediction, limiting the granularity of generated future frames. We will address the limitations in future work.
Appendix A Contributors
@@
Junhao Cai1, Yang Li1, Haoxiang Ma1, Jiangmiao Pang1, Zherui Qiu1, Yang Tian1, Jia Zeng1†, Hongrui Zhu1
@@
Zetao Cai1, Jiafei Cao1, Yilun Chen1, Zeyu He1, Lei Jiang2, Hang Li1, Hengjie Li1, Yufei Liu2, Yanan Lu1, Qi Lv1, Yu Qiao1, Yanqing Shen1, Xu Shi1, Bolun Wang1, Hanqing Wang1, Jiaheng Wang1, Tai Wang1, Xueyuan Wei1, Chao Wu1, Yiman Xie1, Boyang Xing2, Yuqiang Yang1, Yuyin Yang1, Qiaojun Yu1, Feng Yuan1, Jingjing Zhang1, Shenghan Zhang1, Shi Zhang1, Zhuoma Zhaxi1, Bowen Zhou1, Yuanzhen Zhou1, Yunsong Zhou1, Yangkun Zhu1, Yuchen Zhu2